# Spin flip locking by the tunneling and relaxation in a driven double quantum dot with spin-orbit coupling

### v1

@@ III.3 Evolution on the second subharmonic
The reason is that the driving field on the subharmonics operates on lower frequencies and, as a result, the spin flip is achieved on longer times, being stronger affected by the relaxation. The driving frequency for the main harmonic of the $E_{1}$ $E_{4}$ resonance for the considered Zeeman coupling and detuning is in the $18$ $GHz$-range which may not be easily accessible. Hence, the driving on one of the the subharmonics (25) can be a viable option. We will consider the closest subharmonic $s=2$. In Fig.7 we show the spin and coordinate evolution for the second subharmonic $s=2$ on the timescale of $\sim 250$ $ns$ covering $2000$ $T$. All other parameters are the same as in Fig.4. As for the main harmonic, one can see that on the subharmonic the relaxation leads to the dampening of the spin and coordinate oscillations and to the stabilization of the flipped spin. A novel feature is that the stabilized flipped spin projection is somewhat higher for the cases with the relaxation (blue and red curves) compared to the respective coherent oscillations (black curve). We attribute this difference to the progressing effects of the charge and spin relaxation which come into play more quickly (measured in the driving field period) on the subharmonics due to the lower driving frequency, leading to the locking of the electron in the lowest state $E_{2}$ in the right QD with the flipped spin, compared to the state $E_{4}$ in the left QD. The predicted spin flip locking can be considered as an unexpectedly positive effect of the relaxation, maintaining the spin in the flipped state for sufficiently long time.
@@ III.4 Spin flip amplitude dependence on the parameters
The predicted phenomenon of the locked flipped spin at the end of the evolution deserves a more detailed discussion. Namely, it is of interest to gather more information on the locked flipped spin amplitude as a function of the system parameters. First, we look at its dependence on the relaxation times which may vary depending on the structure type and quality. As before, we take equal energy and phase relaxation times for each qubit subsystem, $T_{\varphi}^{(2)}=T_{e}^{(2)}\equiv T^{(2)}$, $T_{\varphi}^{(1)}=T_{e}^{(1)}\equiv T^{(1)}$. In Fig.8 we plot the stabilized (i.e. obtained after a high number of driving periods value of the spin projection $\langle\sigma_{z}\rangle_{t\to\infty}$ for different charge relaxation times $T^{(2)}$ labeled as differently shaped curves and covering a wide interval $(0.01\ldots 0.15)$ $\mu s$ which is consistent with the predictions for the GaAs-based double dots with the detuning Krzywda2021 . Each curve is plotted vs the ratio $T^{(2)}/T^{(1)}$ of the charge and spin relaxation time. All other parameters are the same as in Fig.4. This ratio is always lower than $1$ since the spin relaxation in our system is significantly slower than the charge relaxation. The other parameters are for the main harmonic of the $E_{1}-E_{4}$ resonance shown in Fig.4 and in Fig.6.
Figure 8: Stabilized value of the spin projection $\langle\sigma_{z}\rangle_{t\to\infty}$ for different charge relaxation times $T^{(2)}$ labeled as differently shaped curves, each of them plotted vs the ratio $T^{(2)}/T^{(1)}$ of the charge and spin relaxation time. The other parameters are for the main harmonic of the $E_{1}-E_{4}$ resonance shown in Fig.4 and in Fig.6. The high amplitude of the flipped spin $\sigma_{z}^{\rm{max}}\sim 0.99$ is achieved for moderate and high charge relaxation times $T^{(2)}>0.03\mu s$ and for the long spin relaxation time $T^{(1)}$ when $T^{(2)}/T^{(1)}<0.05$.
@@
Another dependence of the final amplitude of the flipped spin is shown in Fig.9 as a function of the driving strength $V_{d}$. Different curves correspond to different charge relaxation times $T^{(2)}$ with the fixed ratio $T^{(2)}/T^{(1)}=0.1$ which is the typical value in the middle of Fig.8 from which all other parameters are kept. According to Fig.8, for such ratio $T^{(2)}/T^{(1)}$ the best achievable amplitude of $\langle\sigma_{z}\rangle_{t\to\infty}$ is about $0.8$. It is evident from Fig.9 that it can be achieved for a wide range of the charge relaxation time $T^{(2)}$ where all of the curves in the right side of Fig.9 merge together.
@@
Figure 10: Map of the spin flip stabilization time $t_{s}/T$ in units of the driving period $T=0.05$ $ns$ shown vs the spin relaxation time $T^{(1)}$ at fixed ratio $T^{(2)}/T^{(1)}=0.1$ and vs the driving amplitude $V_{d}$. All other parameters are the same as in Fig.9. For a long spin relaxation time (right part of the map) the dependence of $t_{s}/T$ on $V_{d}$ is weak and is determined mainly by the relaxation time itself. At shorter relaxation time (left part of the map) the dependence on $V_{d}$ is sharp, and the stabilized spin flip can be reached much faster.
@@
Our final goal is to explore the dependence on the system parameters for the time when the stabilized spin flip amplitude is reached. In other words, we are interested to know not only how deep is the flip for the locked spin but also how fast this steady state can be reached. In Fig.10 we show the map of the spin flip stabilization time $t_{s}$ in units of the driving period, shown vs the spin relaxation time $T^{(1)}$ and the driving amplitude $V_{d}$. We believe that these two parameters are the most crucial for determining the $t_{s}$. All other parameters are the same as in Fig.9 including the ratio $T^{(2)}/T^{(1)}=0.1$. Each point corresponds to the evolution tracked on a high enough number of the driving field periods $T=0.05$ $ns$ until the stabilized spin projection is achieved. One can see that for the slow relaxation corresponding to the long relaxation time at the right part of the map, the dependence of $t_{s}/T$ is rather shallow, and the obtained times are quite long. This can be explained as a long and coherent-like evolution in the case of slow relaxation where many spin oscillations are visible until the transition to the steady state is manifested, as it can be seen for the examples shown in Fig.7 and Fig.4 by the black and blue curves. As for the area of the faster relaxation in Fig.10, at the left part of the map, the dependence on the driving amplitude $V_{d}$ is much stronger, and the obtained stabilization time becomes significantly shorter if $V_{d}$ is increased. This is due to the faster damping of the coherent spin oscillations which dissipate quickly, unmasking the strong dependence of the spin flip time (the inverse of the Rabi frequency) on the driving field amplitude. To summarize, for the fixed ratio $T^{(2)}/T^{(1)}$ we may interpolate the dependence of the $t_{s}$ in Fig.10 by a simple formula $t_{s}\sim 1/\Omega_{R}+T^{(2)}$. It can be easily verified that this approximation describes a correct topology of the contour plot curves in Fig.10 if we remember that for the exact resonance the Rabi frequency $\Omega_{R}$ is proportional to the driving amplitude $V_{d}$. Another interesting feature of the results in Fig.10 is that an increase in relaxation strength for the charge subsystem, i.e., the decreasing $T^{(2)}$ at a fixed ratio $T^{(2)}/T^{(1)}=0.1$ for a long-lives spin state, can play a positive role here, providing the stabilized spin flip to be reached faster. The stabilized spin flip state can be reached on the interval of several hundreds driving periods which for the parameters in Fig.10 indicates the operation times of $15\ldots 60$ $ns$ which is suitable for the information processing.
@@ IV Conclusions
We have studied the effects of relaxation on the spin and charge evolution in a driven double quantum dot modeling the spin and charge qubits interacting via the spin-orbit coupling. The relaxation effects have been included via the density matrix formalism. The numerical analysis has demonstrated that the relaxation, leading to the trivial damping of the spin oscillations at the electric dipole spin resonance and its second subharmonic, plays a spin flip locking role for the multilevel dynamics where the spin flip is accompanied by an effective tunneling. The spin flip is found to be locked both on the main and on the neighboring subharmonic of the resonance, leading to the prolonged and stable spin flip in the desired quantum dot. The amplitude of the stabilized spin flip can reach as high as $0.99$ in the presence of spin-orbit coupling for a wide range of the relaxation parameters. The stabilized spin flip state can be reached on the interval of about $15\ldots 60$ $ns$, making it suitable for the information processing. The predicted effect can be useful for the spin manipulation techniques in nanostructures and for the new coupled qubit proposals.

### v2

@@ III.3 Evolution on the second subharmonic
The reason is that the driving field on the subharmonics operates on lower frequencies and, as a result, the spin flip is achieved on longer times, being stronger affected by the relaxation. The driving frequency for the main harmonic of the $E_{1}$ $E_{4}$ resonance for the considered Zeeman coupling and detuning is in the $18$ $GHz$-range which may not be easily accessible. Hence, the driving on one of the the subharmonics (25) can be a viable option. We will consider the closest subharmonic $s=2$. In Fig.6 we show the spin and coordinate evolution for the second subharmonic $s=2$ on the timescale of $\sim 250~{}ns$ covering $2000$ $T$. All other parameters are the same as in Fig.4. As for the main harmonic, one can see that on the subharmonic the relaxation leads to the dampening of the spin and coordinate oscillations and to the stabilization of the flipped spin. An interesting and potentially useful feature is that the stabilized flipped spin projection is somewhat higher for the case with the stronger relaxation (red curve vs the blue curve). We attribute this difference to the progressing effects of the charge and spin relaxation which come into play more quickly (measured in the driving field period) on the subharmonics due to the lower driving frequency, leading to the locking of the electron in the lower state $E_{2}$ in the right QD with the flipped spin, compared to the higher state $E_{4}$ in the left QD. The predicted spin flip locking can be considered as an unexpectedly positive effect of the relaxation, maintaining the spin in the flipped state for sufficiently long time.
@@ III.4 Spin flip amplitude dependence on the parameters
The predicted phenomenon of the locked flipped spin at the end of the evolution deserves a more detailed discussion. Namely, it is of interest to gather more information on the locked flipped spin amplitude as a function of the system parameters. First, we look at its dependence on the relaxation times which may vary depending on the structure type and quality. As before, we take equal energy and phase relaxation times for each subsystem, $T_{\varphi}^{(2)}=T_{e}^{(2)}\equiv T^{(2)}$, $T_{\varphi}^{(1)}=T_{e}^{(1)}\equiv T^{(1)}$. In Fig.7 we plot the stabilized (i.e. obtained after a high number of driving periods) value of the spin projection $\langle\sigma_{z}\rangle_{t\to\infty}$ for different charge relaxation times $T^{(2)}$ labeled as differently shaped curves and covering a wide interval $(0.01\ldots 0.15)~{}\mu s$ which is consistent with the predictions for the GaAs-based double dots with the detuning Krzywda2021 . Each curve is plotted vs the ratio $T^{(2)}/T^{(1)}$ of the charge and spin relaxation time. All other parameters are the same as in Fig.4. This ratio is always lower than $1$ since the spin relaxation in our system is significantly slower than the charge relaxation. The other parameters are for the main harmonic of the $E_{1}-E_{4}$ resonance shown in Fig.4 and in Fig.5.
Figure 7: Stabilized value of the spin projection $\langle\sigma_{z}\rangle_{t\to\infty}$ for different charge relaxation times $T^{(2)}$ labeled as differently shaped curves, each of them plotted vs the ratio $T^{(2)}/T^{(1)}$ of the charge and spin relaxation time. The other parameters are for the main harmonic of the $E_{1}-E_{4}$ resonance shown in Fig.4 and in Fig.5. The high amplitude of the flipped spin $\sigma_{z}^{\rm{max}}\sim~{}0.99$ is achieved for moderate and high charge relaxation times $T^{(2)}>0.03\mu s$ and for the long spin relaxation time $T^{(1)}$ when $T^{(2)}/T^{(1)}<0.05$.
@@
Another dependence of the final amplitude of the flipped spin is shown in Fig.8 as a function of the driving strength $V_{d}$. Different curves correspond to different charge relaxation times $T^{(2)}$ with the fixed ratio $T^{(2)}/T^{(1)}=0.1$ which is a typical value in the middle of Fig.7 from which all other parameters are kept. According to Fig.7, for such ratio $T^{(2)}/T^{(1)}$ the best achievable amplitude of $\langle\sigma_{z}\rangle_{t\to\infty}$ is about $0.8$. It is evident from Fig.8 that it can be achieved for a wide range of the charge relaxation time $T^{(2)}$ where all of the curves in the right side of Fig.8 merge together.
@@
Figure 9: Map of the spin flip stabilization time $t_{s}/T$ in units of the driving period $T=0.05$ $ns$ shown vs the charge relaxation time $T^{(2)}$ at fixed ratio $T^{(2)}/T^{(1)}=0.1$ and vs the driving amplitude $V_{d}$. All other parameters are the same as in Fig.8. For a long charge relaxation time (right part of the map) the dependence of $t_{s}/T$ on $V_{d}$ is weak and is determined mainly by the relaxation time itself. At shorter relaxation time (left part of the map) the dependence on $V_{d}$ is sharp, and the stabilized spin flip can be reached much faster.
@@
Our final goal is to explore the dependence on the system parameters for the time when the stabilized spin flip amplitude is reached. In other words, we are interested to know not only how deep is the flip for the locked spin but also how fast this steady state can be reached. In Fig.9 we show the map of the spin flip stabilization time $t_{s}$ defined as the characteristic time of the exponential fit $\sim(1-\exp(-t/t_{s}))$ for the envelope function fitted to the numerically obtained level population $|C_{2}(t)|^{2}$ in Fig.5. The time $t_{s}$ is measured in units of the driving period and is shown in Fig.9 vs the charge relaxation time $T^{(2)}$ and the driving amplitude $V_{d}$. We believe that these two parameters are the most crucial for determining the $t_{s}$. All other parameters are the same as in Fig.8 including the ratio $T^{(2)}/T^{(1)}=0.1$. Each point corresponds to the evolution tracked on a high enough number of the driving field periods $T=0.05$ $ns$ until the stabilized spin projection is achieved. One can see that for the slow relaxation corresponding to the long relaxation time at the right part of the map, the dependence of $t_{s}/T$ is rather shallow, and the obtained times are quite long. This can be explained as a long and coherent-like evolution in the case of slow relaxation where many spin oscillations are visible until the transition to the steady state is manifested, as it can be seen for the examples shown in Fig.6 and Fig.4 by the black and blue curves. As for the area of the faster relaxation in Fig.9, at the left part of the map, the dependence on the driving amplitude $V_{d}$ is much stronger, and the obtained stabilization time becomes significantly shorter if $V_{d}$ is increased. This is due to the faster damping of the coherent spin oscillations which dissipate quickly, unmasking the strong dependence of the spin flip time (the inverse of the Rabi frequency) on the driving field amplitude. To summarize, for the fixed ratio $T^{(2)}/T^{(1)}$ we may interpolate the dependence of the $t_{s}$ in Fig.9 by a simple formula $t_{s}\approx a_{1}/\Omega_{R}+a_{2}T^{(2)}$ where $a_{1,2}$ are the fitting parameters. It can be easily verified that this approximation describes a correct topology of the contour plot curves in Fig.9 if we remember that for the exact resonance the Rabi frequency $\Omega_{R}$ is proportional to the driving amplitude $V_{d}$. Another interesting feature of the results in Fig.9 is that an increase in relaxation strength for the charge subsystem, i.e., the decreasing $T^{(2)}$ at a fixed ratio $T^{(2)}/T^{(1)}=0.1$ for a long-lived spin state, can play a positive role here, providing the stabilized spin flip to be reached faster. The stabilized spin flip state can be reached on the interval of several hundreds driving periods which for the parameters in Fig.9 indicates the operation times of $15\ldots 60$ $ns$ which is suitable for the possible information processing protocols.
@@ IV Conclusions
We have studied the effects of relaxation on the spin and charge evolution in a driven double quantum dot modeling the spin and charge subsystems interacting via the spin-orbit coupling. The relaxation effects have been included via the density matrix formalism. The numerical analysis has demonstrated that the relaxation, leading to the trivial damping of the spin oscillations at the electric dipole spin resonance and its second subharmonic, plays a spin flip locking role for the multilevel dynamics where the spin flip is accompanied by an effective tunneling. The spin flip is found to be locked both on the main and on the neighboring subharmonic of the resonance, leading to the prolonged and stable spin flip in the desired quantum dot. The amplitude of the stabilized spin flip can reach as high as $0.99$ in the presence of spin-orbit coupling for a wide range of the relaxation parameters. The stabilized spin flip state can be reached on the interval of about $15\ldots 60$ $ns$, making it suitable for the information processing. The predicted effect can be useful for the spin manipulation techniques in nanostructures and for the new coupled qubit proposals.